OpenStudy (anonymous):

help simplify -4-(2-6p)

OpenStudy (anonymous):

6p-6

OpenStudy (anonymous):

how did you do that

OpenStudy (anonymous):

you have to make the equation equal 0 so -4(2+6p)=0 1) move 2-6p>> -4=2+6p 2)add 4 to both sides 6+6p 3) the variable always has to be first>>6p+6 so you end up with 6p+6

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

yes, how would I simplify something like this -4r-6-2+r+11

OpenStudy (anonymous):

you have to get all the like terms together

OpenStudy (anonymous):

so, 11,6,and 2 are alike and -4r and r are alike perform the operation and solve...

OpenStudy (anonymous):

so you should have 11-6-2 and -4r+r...what does that give you?